# Charming Darwin: the evolution of QCD parameters across different species

### v1

@@ 4.2 Theory challenges
Full 8.49 4.47 0.30 6.72 5.28
Stat. 4.99 0.85 0.06 3.34 1.59
MC Stat. 3.14 1.42 0.07 2.00 2.71
$\epsilon_{\mathrm{track.}}$ 2.28 2.91 0.20 3.56 1.07
$\sigma_{\mathrm{track.}}$ 0.17 0.03 0.00 0.12 0.06
$\epsilon_{K_{S}^{0}}$ 3.24 1.66 0.15 2.08 1.92
$\sigma_{K_{S}^{0}}$ 0.07 0.06 0.00 0.05 0.04
$\epsilon_{\gamma}$ 1.53 0.93 0.04 0.95 0.94
$\sigma_{\gamma}$ 0.51 0.25 0.04 0.28 1.10
PID 0.03 0.03 0.00 0.04 0.02
$\mathcal{B}(D^{+}\to\eta^{\prime}\ell\nu_{\ell})$ 0.40 0.07 0.02 0.52 1.05
$\mathcal{B}(D^{+}\to\eta\ell\nu_{\ell})$ 1.40 0.32 0.05 1.39 2.17
$\mathcal{B}(D^{+}\to K^{*-}\ell\nu_{\ell})$ 2.63 0.61 0.04 1.68 0.49
$\mathcal{B}(D^{+}\to K^{-}\ell\nu_{\ell})$ 2.18 0.86 0.06 2.24 0.47
$\mathcal{B}(D^{+}\to\omega\ell\nu_{\ell})$ 0.60 0.89 0.08 0.16 0.93
$\mathcal{B}(D^{+}\to\pi^{-}\ell\nu_{\ell})$ 1.60 1.28 0.04 0.52 2.34
$\mathcal{B}(D^{+}\to\rho^{-}\ell\nu_{\ell})$ 0.95 1.05 0.06 1.33 0.65
$\mathcal{B}(D^{+}\to(K\pi)_{S-\mathrm{wave}}\ell\nu_{\ell})$ 4.11 1.75 0.04 1.47 1.32
@@
Full 94.54 101.15 3.05 92.91 94.44
Stat. 34.58 9.63 0.64 28.61 18.03
MC Stat. 25.55 25.78 1.06 21.32 15.39
$\epsilon_{\mathrm{track.}}$ 22.50 28.12 0.96 27.57 21.09
$\sigma_{\mathrm{track.}}$ 8.43 10.08 0.53 5.64 4.20
$\epsilon_{K_{S}^{0}}$ 7.98 6.90 0.22 6.68 12.05
$\sigma_{K_{S}^{0}}$ 8.56 10.05 0.52 5.69 4.22
$\epsilon_{\gamma}$ 7.90 4.28 0.24 11.25 2.37
$\sigma_{\gamma}$ 10.19 7.87 0.56 8.82 7.07
PID 8.56 10.03 0.52 5.68 4.21
$\mathcal{B}(D_{s}\to\eta^{\prime}\ell\nu_{\ell})$ 54.18 9.78 0.21 27.81 10.61
$\mathcal{B}(D_{s}\to\eta\ell\nu_{\ell})$ 33.33 90.26 2.34 60.96 86.46
$\mathcal{B}(D_{s}\to f_{0}\ell\nu_{\ell})$ 14.43 3.31 0.16 8.57 3.12
$\mathcal{B}(D_{s}\to K^{*-}\ell\nu_{\ell})$ 30.21 21.02 0.96 30.77 5.91
$\mathcal{B}(D_{s}\to K^{-}\ell\nu_{\ell})$ 34.83 12.75 0.21 26.67 15.53
$\mathcal{B}(D_{s}\ \to\phi\ell\nu_{\ell})$ 33.44 19.63 0.22 36.21 23.75
@@
Full 50.76 2.79 86.11
Stat. 37.04 1.43 46.08
MC Stat. 27.61 0.86 24.45
$\epsilon_{\mathrm{track.}}$ 7.19 0.55 17.38
$\sigma_{\mathrm{track.}}$ 7.41 0.31 10.01
$\epsilon_{K_{S}^{0}}$ 7.21 0.31 9.85
$\sigma_{K_{S}^{0}}$ 6.92 0.31 9.75
$\epsilon_{\gamma}$ 4.42 0.19 6.28
$\sigma_{\gamma}$ 6.04 0.15 4.85
PID 7.21 0.31 9.85
$\mathcal{B}(\Lambda_{c}\to\Delta^{0}\overline{K}^{0}\ell\nu_{\ell})$ 18.63 1.51 46.83
$\mathcal{B}(\Lambda_{c}\to\Lambda^{0}\ell\nu_{\ell})$ 10.29 0.67 17.19
$\mathcal{B}(\Lambda_{c}\to n^{0}\ell\nu_{\ell})$ 15.96 1.55 51.29
$\mathcal{B}(\Lambda_{c}\to\Sigma^{0}\pi^{0}\ell\nu_{\ell})$ 7.65 0.44 17.09
$\mathcal{B}(\Lambda_{c}\to p^{+}K^{-}\ell\nu_{\ell})$ 15.28 1.42 40.80
@@ 6 Outlook
Comparing the HQE parameters among the different charm species will provide the first test of $SU(3)$ within the HQE and allow valence and non-valence weak annihilation parameters to be separated. Whether or not a HQE analysis of the experimental moments converges will provide an additional and valuable experimental test of the validity of the HQE for charm. Moreover, comparing the HQE parameters between beauty and charm allows for a further test of non-perturbative QCD among different generations. In the infinite mass limit, the HQE parameters in beauty in charm are equal, but in the standard HQE approach they depend in a non-trivial way on the mass of the initial decaying hadron. In Refs. Fael:2019umf ; Gambino:2010jz , the difference between beauty and charm elements is estimated. With the outlined program, these relations can be tested with data.

### v3

@@ 5 Towards an extraction of the Darwin and other HQE operators
Full 8.5 4.5 0.3 6.7 5.3
Stat. 5.0 0.8 0.1 3.3 1.6
MC Stat. 3.1 1.4 0.1 2.0 2.7
$\epsilon_{\mathrm{track.}}$ 2.3 2.9 0.2 3.6 1.1
$\sigma_{\mathrm{track.}}$ 0.2 0.0 0.0 0.1 0.1
$\epsilon_{K_{S}^{0}}$ 3.2 1.7 0.1 2.1 1.9
$\sigma_{K_{S}^{0}}$ 0.1 0.1 0.0 0.1 0.0
$\epsilon_{\gamma}$ 1.5 0.9 0.0 0.9 0.9
$\sigma_{\gamma}$ 0.5 0.2 0.0 0.3 1.1
PID 0.0 0.0 0.0 0.0 0.0
$\mathcal{B}(D^{+}\to\eta^{\prime}\ell\nu_{\ell})$ 0.4 0.1 0.0 0.5 1.1
$\mathcal{B}(D^{+}\to\eta\ell\nu_{\ell})$ 1.4 0.3 0.1 1.4 2.2
$\mathcal{B}(D^{+}\to K^{*-}\ell\nu_{\ell})$ 2.6 0.6 0.0 1.7 0.5
$\mathcal{B}(D^{+}\to K^{-}\ell\nu_{\ell})$ 2.2 0.9 0.1 2.2 0.5
$\mathcal{B}(D^{+}\to\omega\ell\nu_{\ell})$ 0.6 0.9 0.1 0.2 0.9
$\mathcal{B}(D^{+}\to\pi^{-}\ell\nu_{\ell})$ 1.6 1.3 0.0 0.5 2.3
$\mathcal{B}(D^{+}\to\rho^{-}\ell\nu_{\ell})$ 0.9 1.1 0.1 1.3 0.7
$\mathcal{B}(D^{+}\to(K\pi)_{S-\mathrm{wave}}\ell\nu_{\ell})$ 4.1 1.8 0.0 1.5 1.3
@@
Full 94.5 101.2 3.0 92.9 94.4
Stat. 34.6 9.6 0.6 28.6 18.0
MC Stat. 25.6 25.8 1.1 21.3 15.4
$\epsilon_{\mathrm{track.}}$ 22.5 28.1 1.0 27.6 21.1
$\sigma_{\mathrm{track.}}$ 8.4 10.1 0.5 5.6 4.2
$\epsilon_{K_{S}^{0}}$ 8.0 6.9 0.2 6.7 12.1
$\sigma_{K_{S}^{0}}$ 8.6 10.1 0.5 5.7 4.2
$\epsilon_{\gamma}$ 7.9 4.3 0.2 11.2 2.4
$\sigma_{\gamma}$ 10.2 7.9 0.6 8.8 7.1
PID 8.6 10.0 0.5 5.7 4.2
$\mathcal{B}(D_{s}\to\eta^{\prime}\ell\nu_{\ell})$ 54.2 9.8 0.2 27.8 10.6
$\mathcal{B}(D_{s}\to\eta\ell\nu_{\ell})$ 33.3 90.3 2.3 61.0 86.5
$\mathcal{B}(D_{s}\to f_{0}\ell\nu_{\ell})$ 14.4 3.3 0.2 8.6 3.1
$\mathcal{B}(D_{s}\to K^{*-}\ell\nu_{\ell})$ 30.2 21.0 1.0 30.8 5.9
$\mathcal{B}(D_{s}\to K^{-}\ell\nu_{\ell})$ 34.8 12.8 0.2 26.7 15.5
$\mathcal{B}(D_{s}\ \to\phi\ell\nu_{\ell})$ 33.4 19.6 0.2 36.2 23.8
@@
Full 50.8 2.8 86.1
Stat. 37.0 1.4 46.1
MC Stat. 27.6 0.9 24.4
$\epsilon_{\mathrm{track.}}$ 7.2 0.6 17.4
$\sigma_{\mathrm{track.}}$ 7.4 0.3 10.0
$\epsilon_{K_{S}^{0}}$ 7.2 0.3 9.8
$\sigma_{K_{S}^{0}}$ 6.9 0.3 9.8
$\epsilon_{\gamma}$ 4.4 0.2 6.3
$\sigma_{\gamma}$ 6.0 0.1 4.8
PID 7.2 0.3 9.8
$\mathcal{B}(\Lambda_{c}\to\Delta^{0}\overline{K}^{0}\ell\nu_{\ell})$ 18.6 1.5 46.8
$\mathcal{B}(\Lambda_{c}\to\Lambda^{0}\ell\nu_{\ell})$ 10.3 0.7 17.2
$\mathcal{B}(\Lambda_{c}\to n^{0}\ell\nu_{\ell})$ 16.0 1.6 51.3
$\mathcal{B}(\Lambda_{c}\to\Sigma^{0}\pi^{0}\ell\nu_{\ell})$ 7.7 0.4 17.1
$\mathcal{B}(\Lambda_{c}\to p^{+}K^{-}\ell\nu_{\ell})$ 15.3 1.4 40.8
@@ 6 Outlook
Comparing the HQE parameters among the different charm species will provide the first test of $SU(3)$ flavour symmetry within the HQE and allow valence and non-valence weak annihilation parameters to be separated. Whether or not a HQE analysis of the experimental moments converges will provide an additional and valuable experimental test of the validity of the HQE for charm. Moreover, comparing the HQE parameters between beauty and charm allows for a further test of non-perturbative QCD among different generations. In the infinite mass limit, the HQE parameters in beauty in charm are equal, but in the standard HQE approach they depend in a non-trivial way on the mass of the initial decaying hadron. In Refs. Fael:2019umf ; Gambino:2010jz , the difference between beauty and charm elements is estimated. With the outlined program, these relations can be tested with data.